OpenStudy (anonymous):

Which number would you add to three and four-tenths to give an approximate sum of 5?

OpenStudy (anonymous):

5-3 4/10 = x x= 8/5

hero (hero):

Hint: \(x + 3\frac{4}{10} \approx 5\)

hero (hero):

I'm not a big fan of improper fractions

OpenStudy (anonymous):

ME EITHER

hero (hero):

I can show you how to solve these without using improper fractions

OpenStudy (anonymous):

OKAY

hero (hero):

Well basically, we figured out that we need to solve for x. All we have to do is subtract this: \(5 - 3\frac{4}{10}\)

hero (hero):

What we do is separate the whole numbers from the fractions to get this: \((4 - 3)(\frac{10}{10} - \frac{4}{10}\))

hero (hero):

I borrowed 1 from 5 and changed that 1 to 10/10 so that I would have a fraction to subtract from 4/10

hero (hero):

Now I can subtract: \((4 - 3)(\frac{10}{10} - \frac{4}{10})\) becomes \(1\frac{6}{10}\) and that reduces to \(1\frac{3}{5}\) Thus I have done it without improper fractions.

OpenStudy (anonymous):

WOW..

hero (hero):

You can basically use this method with any mixed fractions that need to be added or subtracted. I haven't figured it out a method for multiplying mixed fractions, but I'm close.

hero (hero):

Does anyone have any questions or doubts about my method? Feel free to ask.